# ARM: Refining Multivariate Forecasting with Adaptive Temporal-Contextual Learning

### v1

@@ A.3 Additonal Ablation Studies
MKLS w/o attention 0.129 0.227 0.158 0.255 0.294 0.347 0.367 0.369
MKLS kernel $s=[49]$ 0.129 0.226 0.163 0.260 0.287 0.347 0.367 0.369
[…]
@@ A.3.2 Sub-modules of MKLS
[…]
A.3.3 Other Combinations of ARM
@@
A.3.4 Effect of AUEL Distribution Learning
@@
A.3.5 Effect of AUEL Temporal Pattern Learning: Parameter Size Reduction
In this section, we demonstrate the efficacy of the ”learning temporal patterns” component in AUEL in enhancing model parameter efficiency and reducing the optimal parameter size required in the main predictor, as shown in Table 5. Owing to AUEL’s ability to disentangle univariate effects, the majority of parameters in the main predictor are primarily used for modeling inter-series dependencies, resulting in a significant reduction in the number of parameters needed. From the results, it is evident that on both the larger-scale Electricity dataset and the smaller ETTm1 dataset, the introduction of MoE significantly reduces the optimal model dimension $d$.
@@ A.3.5 Effect of AUEL Temporal Pattern Learning: Parameter Size Reduction
A.4 Model Traning and Hyper-parameters
@@
The ARM (Vanilla) model is trained using the Adam optimizer and MSE loss in Pytorch, with a learning rate of 0.00005 over 100 epochs on each dataset with a early-stopping patience being 30 steps. The first 10% of epochs are for warm-up, followed by a linear decay of learning rate. Owing to ARM’s adaptability to long sequences, unlike baseline models, we only use a lookback window of length 720 for training (and 104 for the ILI dataset). The multi-kernel size $s$ of MKLS is set to [25 145 385]delimited-[]25145385[25\ \ 145\ \ 385][ 25 145 385 ], which is also used as the multi-window size in the AUEL’s adaptive learning of standard deviation.
@@
We run our model on a single Nvidia RTX 3090 GPU. We use a batch size of 32 for most datasets. For some datasets with a larger number of series, if a longer $L_{P}$ renders the running unfeasible, we try to reduce the batch size to 16 or 8, respectively. For the dimension of the main part of the model, we need to adjust it accordingly based on the strength of series-wise relationship. We can mainly consider the scale of a dataset if we are not familiar with the dataset. We employ a Transformer model dimension $d=16$ for most of the small datasets like ETTs, Exchange, ILI, and Multi with less series-wise relationship needed to be learned by our model. For the datasets like Weather, Electricity, and Traffic, which have more series and potentially more series-wise relationship to learn, we raise the dimension $d$ to 64 (you can also refer to Table 5 for more intuition of the setting of $d$. You can basically use these two dimension settings for most of the datasets. As for the number of heads in multi-head attention, we set it to 8. For the number of layers in the Transformer encoder-decoder structure, we use two encoder layers and one decoder layer. We do not apply dropout in the Transformer Encoder; in the MKLS, we set the dropout rate to 0.25; and in the MoE, we set the dropout rate to 0.75. We use a high dropout rate for MoE to avoid overfitting because we set a relatively large $4(L_{I}+L_{P})$ as the hidden dimension of the MoE predictors. As for the number of experts in the MoE module, we use 2 experts for the small datasets stated above and 4 experts for the larger datasets. We set the random seed for the main experiments as 2024.
@@
A.5 Additional Implementation Details of Modules
A.5.1 Additional Implementation Details of AUEL
@@
A.5.2 Additional Implementation Details of MKLS and the Transfering of MKLS to Univariate Models
@@
In ARM (Vanilla), MKLS is applied to a latent model dimension with $d$ channels. For multivariate LTSF Transformers like Autoformer and Informer, they project input data to this latent model dimension, making their MKLS application consistent with Vanilla. However, for univariate models like DLinear and PatchTST, $X^{(t)}$ is not projected onto the latent dimension, but maintaining the original $C$ channels of input series when processing the input series. This inconsistency might skew results in model comparisons, impeding the full potential of MKLS in univariate models. Consequently, for these univariate models, we devised an mixed input and output processing method. After AUEL preprocessing, we project the $C$ series to a dimension of $d-C$ and concatenate it with these series to get a Pre-MKLS input of dimension $d$. Thus, within this $d$-dimensional input, channels from the original series ($C$) coexist with channels representing a mix of multiple series ($d-C$). After receiving the predictor’s output, it is fed into Post-MKLS, then projected back to dimension $C$, merged with the previous result, and subjected to AUEL inverse processing. This strategy seamlessly integrates MKLS’s capacity for handling inter-series relationships into these univariate models, while preserving the univariate models’ proficiency in processing intra-series information independently, providing a fair comparison between the univariate methods and multivariate methods when implementing MKLS module on them.
A.5.3 Additional Implementation Details of Random Dropping
@@
A.5.4 Other Implementation Details
@@
A.6 Computational Costs
We present the comparison of computational costs between our Vanilla+ARM method and the existing LTSF SOTAs, as shown in Table 6. The results are calculated using the ”ptflops” package. We conduct the experiments with the same input format as the data input in ETTm1 dataset. For existing models, we build them with the best hyper-parameter settings stated in their original papers. We set the token dimension of Vanilla to 64 and $d$ of Vanilla+ARM to 16 in order to emphasize the efficiently using of predictor parameters provided by our ARM architecture: the optimal dimension required by the encoder-decoder is reduced after applying ARM, as stated in previous sections.
Table 6: Comparison of computational costs between LTSF models. We use the data input format of ETTm1 to perform this comparison.
Vanilla+ARM Vanilla+ARM Vanilla Vanilla Autoformer Autoformer Informer Informer PatchTST PatchTST
(FLOPs) (Params) (FLOPs) (Params) (FLOPs) (Params) (FLOPs) (Params) (FLOPs) (Params)
$X_{P}=96$ 426M 7.89M 244M 14.9M 10.9G 15.5M 9.41G 11.3M 5.26G 4.87M
$X_{P}=192$ 515M 10.4M 273M 17.5M 11.6G 15.5M 10.1G 11.3M 5.31G 7.08M
$X_{P}=336$ 664M 14.9M 316M 21.9M 12.7G 15.5M 11.2G 11.3M 5.38G 10.4M
$X_{P}=720$ 1.15G 30.6M 431M 37.8M 15.5G 15.5M 14.0G 11.3M 5.58G 19.2M
A.7 Visualization Analysis
A.7.1 Visualization of the Forecasting for Multi Dataset
@@ A.7.2 Randomness of Training
Figure 6: Randomness of Training. The blue line represents the MSE of the combined best baseline results over these eight datasets. The red line, accompanied by corresponding error bars, depicts the best, worst, and average results of ARM (Vanilla) under five settings with random seeds $\in\{2021,2022,2023,2024,2025\}$. On most datasets, the worse results of ARM varying random seeds can still surpass the previously best results from combined baslines.

### v2

@@ A.4 Additonal Ablation Studies
MKLS w/o attention 0.129 0.227 0.158 0.255 0.294 0.347 0.367 0.389
MKLS kernel $s=[49]$ 0.129 0.226 0.163 0.260 0.287 0.347 0.367 0.389
MKLS kernel s=[49 145]s=[49\ \ 145] 0.129 0.230 0.164 0.262 0.292 0.345 0.366 0.387
MKLS kernel s=[49 145 385 673]s=[49\ \ 145\ \ 385\ \ 673] 0.127 0.224 0.156 0.255 0.297 0.349 0.365 0.389
@@ A.4.2 Sub-modules of MKLS
We further examined optimal kernel size $s$ choices within MKLS. It’s worth noting that for other experiments, we utilized s=[49 145 385]s=[49\ \ 145\ \ 385]. Results indicate that, for $L_{I}=720$, this size selection outperforms other combinations. Especially in datasets like Electricity, with numerous series and longer temporal dependencies, proper use of larger convolutional kernels somewhat improved performance.
A.4.3 Other Combinations of ARM
@@
A.4.4 Effect of AUEL Distribution Learning
@@
A.4.5 Reduction and Robustness Improvement of Parameter Size with ARM
In this section, we demonstrate the efficacy of AUEL in enhancing model parameter efficiency and reducing the optimal parameter size required in the main predictor, as shown in Table 5. Owing to AUEL’s ability to disentangle univariate effects, the majority of parameters in the main predictor are primarily used for modeling inter-series dependencies, resulting in a significant reduction in the number of parameters needed. From the results, it is evident that on both the larger-scale Electricity dataset and the smaller ETTm1 dataset, the introduction of MoE significantly reduces the optimal model dimension $d$.
Also, with the existence of ARM, the model become insensitive to the choice of $d$: after the extraction of univariate effect by AUEL, the model parameters of other parts, whose size mainly controlled by model dimension $d$, will focus on modelling the inter-series dependencies. $d$ typically needs to be tuned to accommodate different inter-series relationship strength of datasets. A smaller $d$ helps avoid overfitting in datasets with weak inter-series relationships, whereas a larger $d$ is necessary for datasets with stronger inter-series dependencies. Incorporating Random Dropping mitigates the need for tuning $d$ across various datasets. Random Dropping can be regarded as an adaptive scaler for $d$, allowing the model to automatically adjust to the strength of these relationships. Consequently, $d$ can be set solely based on the number of input series $C$ (we use $d=16$ for $C<20$ and $d=64$ for $C\geq 20$ for all the experiments), with Random Dropping modulating the learning of inter-series relationships. On datasets Exchange (weak inter-series dependencies) and Multi (strong inter-series dependencies) with both of their $C=8$, the results in Table 6 demonstrate the effectiveness of ”R”: using a consistent $d=16$ that potentially fits stronger inter-series dependencies (like Multi), models with Random Dropping (+ARM) perform equally well on both types of datasets, without the overfitting on Exchange dataset observed in models without Random Dropping (+AM).
@@ A.4.5 Reduction and Robustness Improvement of Parameter Size with ARM
Table 6: The insensitivity of the performance regarding the model dimension $d$ with the existence of Random Dropping. We conducted the experiments with a consistent $d=16$ on two datasets with the same number of series $C=8$ but different strength of inter-series relationship. Exchange dataset have weak inter-series link which inherently requires less parameters $d<16$ to model it, while Multi dataset have stronger inter-series dependencies and fits the scenario of $d=16$. With ARM comparing to AM, models perform equally well on datasets with both strong and weak inter-series dependencies. In other words, random dropping can provide more improvement on the datasets with weaker inter-series dependencies when $d$ is high enough based on $C$.
Models Autoformer+AM Autoformer+ARM Informer+AM Informer+ARM Vanilla+AM Vanilla+ARM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Exchange (96) 0.089 0.21 0.081 0.201 0.090 0.212 0.086 0.207 0.085 0.206 0.078 0.197
Exchange (336) 0.332 0.417 0.298 0.395 0.338 0.420 0.312 0.404 0.304 0.398 0.252 0.367
Average% 100.0% 100.0% 90.0% 95.1% 100.0% 100.0% 93.0% 96.7% 100.0% 100.0% 84.8% 93.4%
Multi (96) 0.037 0.130 0.038 0.130 0.051 0.164 0.051 0.166 0.034 0.129 0.032 0.125
Multi (336) 0.181 0.302 0.175 0.300 0.182 0.305 0.183 0.304 0.172 0.291 0.164 0.286
Average% 100.0% 100.0% 97.7% 99.5% 100.0% 100.0% 100.4% 100.2% 100.0% 100.0% 95.1% 97.9%
A.5 Model Traning and Hyper-parameters
@@
The ARM (Vanilla) model is trained using the Adam optimizer and MSE loss in Pytorch, with a learning rate of 0.00005 over 100 epochs on each dataset with a early-stopping patience being 30 steps. The first 10% of epochs are for warm-up, followed by a linear decay of learning rate. Owing to ARM’s adaptability to long sequences, unlike baseline models, we only use a lookback window of length 720 for training (and 104 for the ILI dataset). The multi-kernel size $s$ of MKLS is set to [25 145 385][25\ \ 145\ \ 385], which is also used as the multi-window size in the AUEL’s adaptive learning of standard deviation.
@@
We run our model on a single Nvidia RTX 3090 GPU. We use a batch size of 32 for most datasets. For some datasets with a larger number of series, if a longer $L_{P}$ renders the running unfeasible, we try to reduce the batch size to 16 or 8, respectively. For the dimension of the main part of the model, we need to adjust it accordingly based on the strength of series-wise relationship. We can mainly consider the scale of a dataset if we are not familiar with the dataset. We employ a Transformer model dimension $d=16$ for most of the small datasets ($C<20$) like ETTs, Exchange, ILI, and Multi with less series-wise relationship needed to be learned by our model. For the datasets like Weather, Electricity, and Traffic ($C\geq 20$), which have more series and potentially more series-wise relationship to learn, we raise the dimension $d$ to 64 (you can also refer to Table 5 for more intuition of the setting of $d$). You can basically use these two dimension settings for most of the datasets. As for the number of heads in multi-head attention, we set it to 8. For the number of layers in the Transformer encoder-decoder structure, we use two encoder layers and one decoder layer. We do not apply dropout in the Transformer Encoder; in the MKLS, we set the dropout rate to 0.25; and in the MoE, we set the dropout rate to 0.75. We use a high dropout rate for MoE to avoid overfitting because we set a relatively large $4(L_{I}+L_{P})$ as the hidden dimension of the MoE predictors. As for the number of experts in the MoE module, we use 2 experts for the small datasets stated above and 4 experts for the larger datasets. We set the random seed for the main experiments as 2024.
@@
A.6 Additional Implementation Details of Modules
A.6.1 Additional Implementation Details of AUEL
@@
A.6.2 Additional Implementation Details of MKLS and the Transfering of MKLS to Univariate Models
@@
In ARM (Vanilla), MKLS is applied to a latent model dimension with $d$ channels. For multivariate LTSF Transformers like Autoformer and Informer, they project input data to this latent model dimension, making their MKLS application consistent with Vanilla. However, for univariate models like DLinear and PatchTST, $X^{(t)}$ is not projected onto the latent dimension, but maintaining the original $C$ channels of input series when processing the input series. This inconsistency might skew results in model comparisons, impeding the full potential of MKLS in univariate models. Consequently, for these univariate models, we devised an mixed input and output processing method. After AUEL preprocessing, we project the $C$ series to a dimension of dCd-C and concatenate it with these series to get a Pre-MKLS input of dimension $d$. Thus, within this $d$-dimensional input, channels from the original series ($C$) coexist with channels representing a mix of multiple series (dCd-C). After receiving the predictor’s output, it is fed into Post-MKLS, then projected back to dimension $C$, merged with the previous result, and subjected to AUEL inverse processing. This strategy seamlessly integrates MKLS’s capacity for handling inter-series relationships into these univariate models, while preserving the univariate models’ proficiency in processing intra-series information independently, providing a fair comparison between the univariate methods and multivariate methods when implementing MKLS module on them.
A.6.3 Additional Implementation Details of Random Dropping
@@
A.6.4 Other Implementation Details
@@
A.7 Computational Costs
We present the comparison of computational costs between our Vanilla+ARM method and the existing LTSF SOTAs, as shown in Table 7. The results are calculated using the ”ptflops” package. We conduct the experiments with the same input format as the data input in ETTm1 dataset. For existing models, we build them with the best hyper-parameter settings stated in their original papers. We set the token dimension of Vanilla to 64 and $d$ of Vanilla+ARM to 16 in order to emphasize the efficiently using of predictor parameters provided by our ARM architecture: the optimal dimension required by the encoder-decoder is reduced after applying ARM, as presented in Table 5.
Table 7: Comparative Analysis of Computational Costs in LTSF Models. This comparison utilizes the ETTm1 data format for constructing model inputs.
Vanilla+ARM Vanilla+ARM Vanilla Vanilla Autoformer Autoformer
(FLOPs) (Params) (FLOPs) (Params) (FLOPs) (Params)
$X_{P}=96$ 426M 7.89M 244M 14.9M 10.9G 15.5M
$X_{P}=192$ 515M 10.4M 273M 17.5M 11.6G 15.5M
$X_{P}=336$ 664M 14.9M 316M 21.9M 12.7G 15.5M
$X_{P}=720$ 1.15G 30.6M 431M 37.8M 15.5G 15.5M
Informer Informer PatchTST PatchTST DLinear DLinear
(FLOPs) (Params) (FLOPs) (Params) (FLOPs) (Params)
$X_{P}=96$ 9.41G 11.3M 5.26G 4.87M 3.06M 138K
$X_{P}=192$ 10.1G 11.3M 5.31G 7.08M 6.10M 277K
$X_{P}=336$ 11.2G 11.3M 5.38G 10.4M 10.7M 485K
$X_{P}=720$ 14.0G 11.3M 5.58G 19.2M 22.8M 1.04M
A.8 Predictability Analysis
To illustrate the variability in predictability across datasets, we conduct a simple analysis in this section, as shown in Figure 5. In each of the eight datasets shown in the figure, we randomly select subseries of length 432, using the first 336 steps as the training set and the final 96 steps as the testing set to fit traditional univariate time series models. The models we employ include naive repeat, ARMA(p, q), Simple Moving Average with different lookback windows (SMA), Exponential Smoothing with different alpha values (ES), Random Forest with different lookback windows (RFR), and MLP with various hidden layer dimensions and lookback windows. We calculate the RMSE on the testing set as the performance metric for each model. To compare the relative performance of different models on various series, we normalize and invert the RMSE for each series, then standardize across the model dimension. This process ensures that a higher value corresponds to better relative model performance. We visualize the transformed performance evaluations of different models on various datasets using boxplots.
As shown in the figure, in datasets where time series with regular patterns dominate, such as Electricity, more complex models like MLP and Random Forest demonstrate superior performance for these regularities, and longer lookback lengths yield better results. In contrast, in datasets with weaker regularity and sudden changes, like Exchange and ILI, methods that either focus on a shorter lookback view or assign greater weight to recent datapoints, such as SMA and ES, perform better. This predictability analysis indicates the necessity of adaptively adjusting the lookback view and complexity of a model for time series with different characteristics. In ARM, we use AUEL, MKLS, and Random Dropping to accommodate these characteristics changes, effectively handling the significant differences observed across various datasets.
Figure 5: Predictability Analysis of Datasets (Elec. means Electricity)
A.9 Visualization Analysis
A.9.1 Visualization of the Forecasting for Multi Dataset
@@ A.9.2 Randomness of Training
Figure 7: Randomness of Training. The blue line represents the MSE of the combined best baseline results over these eight datasets. The red line, accompanied by corresponding error bars, depicts the best, worst, and average results of ARM (Vanilla) under five settings with random seeds $\in{2021,2022,2023,2024,2025}$. On most datasets, the worse results of ARM varying random seeds can still surpass the previously best results from combined baslines.
A.10 Future Works
For future research, potential applications of ARM modules in other time series tasks can be explored. Given that MKLS and Random Dropping decouple relationships between series, it would be worth investigating their viability in time series classification and anomaly detection. Additionally, as AUEL effectively extracts a series’ intrinsic effects, enhancing the extraction of inter-series relationships, its application in tasks like time series clustering can be further explored.